PHOTOCHEMISTRY OF BIACETYL-d_{6} ISOLATED IN INERT GAS MATRICES

JARED VARALY and \underline{C. A. BAUMANN}, Department of Chemistry, University of Scranton, Scranton, PA 18510-4626.

We report the results of ultraviolet photolysis of biacetyl-d_{6} ((CD_{3}CO)_{2}) trapped in solid nitrogen, argon, and krypton. The infrared spectra obtained prior to photolysis are in agreement with the gas phase results obtained by Durig and coworkers ^{a}. The photoproducts were characterized via infrared spectroscopy. Irradiation at 405 nm results in the production of CO, CD_{3}CO, and CD_{3}, which are the photoproducts expected from gas phase biacetyl-h_{6} studies at this wavelength ^{b}. Recombination of the trapped photoproducts leads to the production of other species not observed in the gas phase photolysis, including ketene-d_{2} (CD_{2}CO) and acetaldehyde-d_{4} (CD_{3}CDO). Mechanisms for the production of these and other secondary photoproducts will be discussed.

aJ. R. Durig, S. E. Hannum, S. C. Brown, J. Phys. Chem. 75, 1946 (1971).

bG. F. Sheats, W. A. Noyes, Jr., J. Am. Chem. Soc. 77, 1421 (1955).